# Enhancing Low-light Light Field Images with A Deep Compensation Unfolding Network

### v2

@@ Abstract
This paper presents a novel and interpretable end-to-end learning framework, called the deep compensation unfolding network (DCUNet), for restoring light field (LF) images captured under low-light conditions. DCUNet is designed with a multi-stage architecture that mimics the optimization process of solving an inverse imaging problem in a data-driven fashion. The framework uses the intermediate enhanced result to estimate the illumination map, which is then employed in the unfolding process to produce a new enhanced result. Additionally, DCUNet includes a content-associated deep compensation module at each optimization stage to suppress noise and illumination map estimation errors. To properly mine and leverage the unique characteristics of LF images, this paper proposes a pseudo-explicit feature interaction module that comprehensively exploits redundant information in LF images. The experimental results on both simulated and real datasets demonstrate the superiority of our DCUNet over state-of-the-art methods, both qualitatively and quantitatively. Moreover, DCUNet preserves the essential geometric structure of enhanced LF images much better. The code will be publicly available at https://github.com/lyuxianqiang/LFLL-DCU.
@@ I Introduction
Low-light enhancement for LF images is still preliminary, and only a few methods have been proposed. The existing methods [13, 14, 15] are mainly focused on exploiting redundant information in LF images by improving the spatial-angular separable (SAS) convolution [14] or utilizing predefined view stacks [13, 15]. However, these methods cannot use the LF properties well, thus limiting their performance. For instance, the neighboring SAI stacks used in [13, 15] can only perceive the local information of the LF and inevitably suffer from the aliasing problem and artifacts when processing large disparity LF. Besides, they are purely data-driven, limiting their generalizability to some extent. Moreover, intense noise in low-light LF images will destroy the structural coherence of the LF and reduce the signal-to-noise ratio in extremely dark situations, which puts forward higher requirements for feature extraction.
In this paper, we tackle this challenging problem by proposing a deep compensation unfolding network (DCUNet), guided by the physical model of low-light LF imaging and the constrained optimization process of solving an inverse problem with the half quadratic splitting algorithm [16, 17]. Specifically, DCUNet is constructed with a multi-stage architecture that uses a deep optimization module to predict the enhancement results by leveraging the estimated illumination map and deep compensation for noise and illumination estimation error at each stage. To utilize the unique characteristic of LF images, i.e., the repeated observations (SAIs) of an identical scene, we propose a pseudo-explicit feature interaction module that emphasizes the traits of the epipolar plane images (EPIs) and converges rays from SAIs. Moreover, DCUNet considers the impact of noise in modeling and network design to enhance low-light LF with severe noise.
@@
  • a novel deep unfolding network for enhancing low-light LF images, comprising of illumination map estimation, content-associated deep compensation, and deep optimization;
  • an insight into unfolding-based low-light enhancement, emphasizing the necessity of integrating noise impact within the optimization process; and
@@ II-A Low-light Image/Video Enhancement
With the emergence of deep learning, large amounts of data and deeper networks dramatically benefit low-light image enhancement tasks. Ren et al. [7] suggested a deep hybrid network for improving low-light images that simultaneously uses two streams to learn global information and salient structures. Xu et al. [8] proposed a frequency-based decomposition and enhancement model for low-light image enhancement that recovered picture content at low frequency and details at high-frequency. EnlightenGAN [10], an unsupervised GAN-based approach, learns to improve low-light images from unpaired low/normal light data. Guo et al. proposed a zero-reference curve estimation network (Zero-DCE) [24] and accelerated version (Zero-DCE++) [9] that formulate the light enhancement as a task of image-specific curve estimating, which takes a low-light image as input and generates high-order curves. Ma et al. [25] proposed a self-calibrated illumination learning framework for brightening images in real-world low-light scenarios. It includes a self-calibrated module that realizes the convergence between the results of each stage, producing gains that only use a single basic block for inference, which drastically reduces computation cost. In contrast to existing techniques that rely on pixel-wise reconstruction losses and deterministic processes, Wang et al. [26] proposed a normalizing flow model trained with a negative log-likelihood loss, which takes low-light images or features as conditions. This approach naturally characterizes the structural context and measures the visual distance in the image manifold.
@@ II-B Low-light LF Image Enhancement
Existing light-field low-light enhancement approaches tend to address data containing harsh noise in extremely dark environments. These approaches require enhancing brightness while maintaining view consistency across SAIs. Lamba et al. [13] developed a low-light L3F dataset and a two-stage L3Fnet architecture, which includes a global representation block to encode the LF geometry and a view reconstruction block to restore each view. To restore dark LFs efficiently, they propose a three-stage network [35] that utilizes global, local, and view-specific information in low-light LFs and fuses them using an RNN-inspired feedforward network. Zhang et al. [14] extended the Retinex-Net to LF, developing a learning-based decomposition-enhancement method to decouple the complex task into several sub-tasks. Lamba et al. [36] proposed a three-stage network that utilizes global, local, and view-specific information for fast and efficient low-light restoration in an RNN-inspired network. Zhang et al. [37] utilized the Transformer for efficient low-light LF image restoration. Wang et al. [15] proposed a multiple-stream progressive restoration network for low-light LF enhancement. These algorithms are mostly driven by data, and the network architecture is based on experience, which is not easily interpretable. Additionally, the LF’s benefits for enhancing brightness are not fully utilized.
Figure 1: Flowchart of the proposed end-to-end learning-based framework for low-light LF image enhancement, namely DCUNet. DCUNet mainly consists of three modules: illumination map estimation, content-associated deep compensation, and a deep optimization module.
@@ III Problem Formulation
In low-light situations, the imaging process produces a variety of loud noises, including shot noise, dark current, and readout noise. The readout noise follows a Gaussian distribution, while the shot noise and dark current both follow Poisson distributions [38, 39]. As Poissonian-Gaussian noise can be formulated as an additive noise through heteroskedastic normal approximation [40], we thus formulate the LF imaging process under low-light conditions as
@@
$\begin{split}\min_{\textbf{L}_{n},\bm{\nu},\textbf{N}}&\left\|\textbf{L}_{d}-\textbf{I}\odot\textbf{L}_{n}-\textbf{N}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{\nu})\\ &s.t.\quad\textbf{L}_{n}=\bm{\nu}.\end{split}$ (3)
where $\lambda$ is a positive penalty parameter, and the auxiliary variable $\bm{\nu}$ is used to facilitate optimization [41]. It is worth noting that there are two key differences between this optimization problem and the general optimization process. First, Eq. (3) takes into account the impact of strong noise on the optimization process by including the noise term N in the objective function. Second, the illumination map I is considered as a signal-dependent variable that is related to the optimization result $\textbf{L}_{n}$ rather than a free variable.
Based on the half-quadratic splitting method [16, 17], we can solve the problem in Eq. (3) by alternatively and iteratively tackling the following two sub-problems with the penalty parameter $\mu>0$ for the data likelihood term and the regularization term:
@@
where $k$ is the iteration index. With an estimated illumination map $\textbf{I}^{k}$ and an appropriate deep regularization term, we can obtain the enhancement result by solving Eq. (4).
@@ IV Proposed Method
Figure 3: Demonstration of the utilization of redundant information by different strategies in an epipolar plane image (EPI) perspective. The EPIs shown have a large disparity and severe noise.
@@ IV-A Pseudo-explicit Feature Interaction
Analysis of different feature extraction strategies. Yeung et al. [42] proposed the spatial-angular separable (SAS) convolution, which builds upon the general spatial convolution by using spatial convolution to extract SAI’s features and angular convolution (as shown in Fig. 3(a)) to extract angular features. The proposed SAS convolution associates the information of different SAIs through multiple convolution operations in an indirect way. However, this indirect way of associating angular information may not be effective in handling LF images with large disparity and severe noise. Fig. 3(b) shows the aligned EPI line via the depth-based warping operation, which can be directly fused/averaged to improve brightness and suppress noise. Fig. 3(c) uses the neighboring SAIs of the target SAI to assist feature extraction [13, 15]. However, only local redundant information can be exploited. As depicted in Fig. 3(d), the direct processing of EPI information represents a highly effective approach for addressing large disparity LFs. To achieve comprehensive execution of high-dimensional feature interaction in the absence of depth information, we underline the crucial significance of EPI convolutions, which enable the direct correlation of information across distinct SAIs.
Figure 5: Detail architecture of the proposed pseudo-explicit feature (DPEF) interaction module. The input of the DPEF encompasses dense feature stack (spatial, angular, and epipolar plane image) information, with stack parameter $D\in\left\{1,...,L\right\}$, where $L$ represents the layer number. To balance the quality of the low-light LF enhancement and computational expenses, we ultimately settled on 6 layers, and 32 channel numbers in our framework.
@@ IV-B Illumination Map Estimation
The illumination map plays a crucial role in the optimization process for low-light LF image enhancement. In [30], the illumination map and the enhancement result are optimized simultaneously as independent variables. However, our method takes a different approach by estimating the illumination map using the enhancement results. Specifically, at the $k$-th stage, based on the enhanced result of the preceding stage $\textbf{L}_{n}^{k-1}$ and the input low-light LF image $\textbf{L}_{d}$, we compute a coarse illumination map $\tilde{\textbf{I}}^{k}$, i.e.,
@@ IV-C Content-associated Deep Compensation
The low-light LF images, affected by severe Gaussian and Poisson noise, exhibit a reduced signal-to-noise ratio, imposing higher requirements on the enhancement algorithm. In the context of the optimization process, disregarding the noise N in Eq. (3) introduces a non-negligible bias within the solution space. Incorporating a deep regularization term in such conditions necessitates the regularization term to account for both the intrinsic characteristics of the data and the influences of noise. Compared with regularization terms that solely focus on data characteristics, deep regularization terms that couple multiple data features pose greater challenges in terms of learning. Therefore, it is crucial to consider the content-associated noise N during optimization. In addition to noise, errors in illumination map estimation can also affect optimization results. Therefore, we propose a content-associated deep compensation term $\bm{\delta}$ to attenuate the negative effects of noise and illumination map error on the optimization process.
@@ IV-D Deep Optimization
With the estimated illumination map $I^{k}$ and content-associated deep compensation $\delta^{k}$ in the $k$-th iteration, we address the optimization problem outlined in Eq. (4) through a deep optimization module that comprises a defined computation and a deep regularization term. Specifically, we construct the deep implicit regularization term by utilizing pseudo-explicit feature interaction modules with a dense-skip connection. The solution to Eq. (4) can be expressed as
@@ V-A Experiment Settings
Datasets. Both synthetic LF images from the Inria [48] and HCI [49] benchmarks and real-world LF images captured with a Lytro Illum camera provided by L3F dataset [13] were employed to train and test. Specifically, for the synthetic LF dataset, 22 LF images of dimensions $5\times 5\times 512\times 512$ from HCI and 33 LF images of the same dimensions from Inria were selected to form the training set. The test set contains 2 LF images from HCI and 4 LF images from Inria. For the L3F dataset [13], we use the same training and test set partitioning provided by the dataset, and the dimensions of the LF data are $7\times 7\times 434\times 625$. Note that L3Fnet needs additional SAIs as the input to their view reconstruction block, so the input angular resolution of L3FNet for synthetic LF and real-world LF are $7\times 7$ and $9\times 9$, respectively.
Figure 9: Visual comparisons of enhanced center SAI by different methods on the L3F dataset [13]. Selected regions have been zoomed in for better comparison. It is recommended to view this figure by zooming in.
@@ V-B Comparison with State-of-the-Art Methods
Quantitative comparisons on the synthetic LF dataset. Two strategies were used to simulate the low-light LF images. One is to follow [13] to simulate the data in a fixed illumination environment and uses the same photoelectron amplification factor $\alpha=0.2$, denoted as Syn-F. The other is that the simulation data is in a dynamic illumination environment, and the photoelectron amplification factor changes within a particular range $\alpha\in\left[0.1,0.3\right]$, denoted as Syn-D. The standard variance of the adding zero-mean Gaussian noise is 20 and 15, respectively.
@@
Visual comparisons. Fig. 8 and Fig. 9 show the visual comparisons of enhanced LF images by different methods. It can be observed that our method produces better visual results than all the compared methods under all tasks. Specifically, our method can produce sharper edges at the occlusion boundaries and better high-frequency details at texture regions than other approaches. Moreover, the visual effects and overall brightness of our results are closest to the ground truth. As shown in Fig. 8, Ours and Ours-D can enhance the low-light data with intense noise and obtain clear images.
Figure 10: Visual comparisons of the center depth maps derived by applying the LF depth estimation method in [4] to the enhanced LF images by different methods. Note that the ground-truth depth of this synthetic dataset is available.

### v3

@@ Abstract
This paper presents a novel and interpretable end-to-end learning framework, called the deep compensation unfolding network (DCUNet), for restoring light field (LF) images captured under low-light conditions. DCUNet is designed with a multi-stage architecture that mimics the optimization process of solving an inverse imaging problem in a data-driven fashion. The framework uses the intermediate enhanced result to estimate the illumination map, which is then employed in the unfolding process to produce a new enhanced result. Additionally, DCUNet includes a content-associated deep compensation module at each optimization stage to suppress noise and illumination map estimation errors. To properly mine and leverage the unique characteristics of LF images, this paper proposes a pseudo-explicit feature interaction module that comprehensively exploits redundant information in LF images. The experimental results on both simulated and real datasets demonstrate the superiority of our DCUNet over state-of-the-art methods, both qualitatively and quantitatively. Moreover, DCUNet preserves the essential geometric structure of enhanced LF images much better. The code is publicly available at https://github.com/lyuxianqiang/LFLL-DCU.
@@ I Introduction
Low-light enhancement for LF images is still preliminary, and only a few methods have been proposed. The existing methods [13, 14, 15, 16] are mainly focused on exploiting redundant information in LF images by improving the spatial-angular separable (SAS) convolution [14] or utilizing predefined view stacks [13, 15, 16]. However, these methods cannot use the LF properties well, thus limiting their performance. For instance, the neighboring SAI stacks used in [13, 15] and selected SAIs in [16] can only perceive the local information of the LF and inevitably suffer from the aliasing problem and artifacts when processing large disparity LF. Besides, they are purely data-driven, limiting their generalizability to some extent. Moreover, intense noise in low-light LF images will destroy the structural coherence of the LF and reduce the signal-to-noise ratio in extremely dark situations, which puts forward higher requirements for feature extraction.
In this paper, we tackle this challenging problem by proposing a deep compensation unfolding network (DCUNet), guided by the physical model of low-light LF imaging and the constrained optimization process of solving an inverse problem with the half quadratic splitting algorithm [17, 18]. Specifically, DCUNet is constructed with a multi-stage architecture that uses a deep optimization module to predict the enhancement results by leveraging the estimated illumination map and deep compensation for noise and illumination estimation error at each stage. To utilize the unique characteristic of LF images, i.e., the repeated observations (SAIs) of an identical scene, we propose a pseudo-explicit feature interaction module that emphasizes the traits of the epipolar plane images (EPIs) and converges rays from SAIs. Moreover, DCUNet emphasizes the necessity of integrating noise impact within the optimization process to enhance low-light LF with severe noise.
@@
  • a novel deep unfolding network for enhancing low-light LF images, featured with illumination map estimation, content-associated deep compensation, and deep optimization;
@@ II-A Low-light Image/Video Enhancement
With the emergence of deep learning, large amounts of data and deeper networks dramatically benefit low-light image enhancement tasks. Ren et al. [7] suggested a deep hybrid network for improving low-light images that simultaneously uses two streams to learn global information and salient structures. Xu et al. [8] proposed a frequency-based decomposition and enhancement model for low-light image enhancement that recovered picture content at low frequency and details at high frequency. EnlightenGAN [10], an unsupervised GAN-based approach, learns to improve low-light images from unpaired low/normal light data. Guo et al. proposed a zero-reference curve estimation network (Zero-DCE) [25] and accelerated version (Zero-DCE++) [9] that formulate the light enhancement as a task of image-specific curve estimating, which takes a low-light image as input and generates high-order curves. Ma et al. [26] proposed a self-calibrated illumination learning framework for brightening images in real-world low-light scenarios. Zheng et al. [27] employed a deep unfolding total variation module to estimate the noise level map, which was then used to assist the subsequent image restoration module. In contrast to existing techniques that rely on pixel-wise reconstruction losses and deterministic processes, Wang et al. [28] proposed a normalizing flow model trained with a negative log-likelihood loss, which takes low-light images or features as conditions. This approach naturally characterizes the structural context and measures the visual distance in the image manifold.
@@ II-B Low-light LF Image Enhancement
Existing light-field low-light enhancement approaches tend to address data containing harsh noise in extremely dark environments. These approaches require enhancing brightness while maintaining view consistency across SAIs. Lamba et al. [13] developed a low-light L3F dataset and a two-stage L3Fnet architecture, which includes a global representation block to encode the LF geometry and a view reconstruction block to restore each view. To restore dark LFs efficiently, they propose a three-stage network [37] that utilizes global, local, and view-specific information in low-light LFs and fuses them using an RNN-inspired feedforward network. Zhang et al. [14] extended the Retinex-Net to LF, developing a learning-based decomposition-enhancement method to decouple the complex task into several sub-tasks. Lamba et al. [38] proposed a three-stage network that utilizes global, local, and view-specific information for fast and efficient low-light restoration in an RNN-inspired network. Zhang et al. [39] utilized the Transformer for efficient low-light LF image restoration. Wang et al. [15] proposed a multiple-stream progressive restoration network for low-light LF enhancement. Wang et al. [16] proposed a parallel multi-scale network (PMSNet) that employs an Unet-like architecture to process light field features at different scales. The PMSNet network leverages both local and global information to achieve improved enhancement results, similar to the approach adopted by L3FNet [13]. These algorithms are mostly driven by data, and the network architecture is based on experience, which is not easily interpretable. Additionally, the LF’s benefits for enhancing brightness are not fully utilized.
Figure 2: Flowchart of the proposed end-to-end learning-based framework for low-light LF image enhancement, namely DCUNet. DCUNet consists of three main modules: Illumination Map Estimation (IE), Content-associated Deep Compensation (CDC), and Deep Optimization (DO). The CDC module compensates for the solution bias of the enhanced intermediate results $\textbf{L}_{n}^{k-1}$ and estimated illumination map $\textbf{I}^{k}$. The DO module is designed to mimic the optimization process and iteratively refine the enhanced result.
@@ III Problem Formulation
In low-light situations, the imaging process produces a variety of loud noises, including shot noise, dark current, and readout noise. The readout noise follows a Gaussian distribution, while the shot noise and dark current both follow Poisson distributions [40, 41]. As Poissonian-Gaussian noise can be formulated as an additive signal-dependent noise through heteroskedastic normal approximation [42], we thus formulate the LF imaging process under low-light conditions as
@@
$\begin{split}\min_{\textbf{L}_{n},\bm{\nu},\textbf{I},\textbf{N}}&\left\|\textbf{L}_{d}-\textbf{I}\odot\textbf{L}_{n}-\textbf{N}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{\nu})\\ &s.t.\quad\textbf{L}_{n}=\bm{\nu}.\end{split}$ (3)
where $\lambda$ is a positive penalty parameter, and the auxiliary variable $\bm{\nu}$ is used to facilitate optimization [43]. It is worth noting that there are two key differences between this optimization problem and the general optimization process. First, Eq. (3) takes into account the impact of strong noise on the optimization process by including the noise term N in the objective function. Second, the illumination map I and noise term N are considered as signal-dependent variables that are related to the optimization result $\textbf{L}_{n}$ rather than free variables.
Consequently, we can leverage the $\textbf{L}_{n}^{k-1}$ to calculate the $\textbf{I}^{k}$ and $\textbf{N}^{k}$, and employ the half-quadratic splitting method [17, 18] to solve the problem in Eq. (3) by alternatively and iteratively tackling the following two sub-problems:
@@
where $k$ is the iteration index, $\mu>0$ is the penalty parameter. With the estimated $\textbf{I}^{k}$, $\textbf{N}^{k}$, and a deep regularization term, we can obtain the enhancement result by solving Eq. (4).
@@ IV Proposed Method
Figure 3: Demonstration of the utilization of redundant information by different strategies in an epipolar plane image (EPI) perspective. The orange blocks with different saturations represent corresponding pixels affected by severe noise, and the green bounding boxes show the scope of different operations.
@@ IV-A Pseudo-explicit Feature Interaction
Analysis of different feature extraction strategies. Yeung et al. [44] proposed the spatial-angular separable (SAS) convolution, which builds upon the general spatial convolution by using spatial convolution to extract SAI’s features and angular convolution (as shown in Fig. 3(a)) to extract angular features. The angular convolution associates the information of different SAIs through multiple convolution operations in an indirect way, which may not be effective in handling LF images with large disparity and severe noise. Fig. 3(b) shows the aligned EPI line via the depth-based warping operation, which can be directly fused/averaged to improve brightness and suppress noise. Fig. 3(c) uses the neighboring SAIs of the target SAI to assist feature extraction [13, 15]. However, only local redundant information can be exploited. As depicted in Fig. 3(d), the direct processing of EPI information [45, 46, 47] represents a highly effective approach for addressing large disparity low-light LFs. Therefore, in addition to using EPI convolution, we need to further emphasize the crucial significance of EPI information.
Figure 5: Detail architecture of the proposed pseudo-explicit feature (DPEF) interaction module. The input of the DPEF encompasses dense feature stack (spatial, angular, and epipolar plane image) information, with stack parameter $D\in\left\{1,...,L\right\}$, where $L$ represents the layer number. The kernel size is 3 and the stride is 1 for all the basic 4D-2D convolutions. To balance the quality of the low-light LF enhancement and computational expenses, we ultimately settled on 6 layers, and 32 channel numbers in our framework.
@@ IV-B Illumination Map Estimation
The illumination map plays a crucial role [29, 32, 39] for Retinex-based low-light image enhancement. In the deep unfolding process of [32], the illumination map and the enhancement result are optimized simultaneously as independent variables. However, our method takes a different approach by estimating the illumination map using the enhancement results. Specifically, at the $k$-th stage, based on the enhanced result of the preceding stage $\textbf{L}_{n}^{k-1}$ and the input low-light LF image $\textbf{L}_{d}$, we compute a coarse illumination map $\tilde{\textbf{I}}^{k}$, i.e.,
@@ IV-C Content-associated Deep Compensation
Low-light light field images with low signal-to-noise ratios pose significant challenges to enhancement algorithms [25, 27]. In the context of the deep unfolding process, disregarding the noise term N in Eq. (3) introduces a non-negligible bias in the solution spaces of the intermediate results $L_{n}^{k-1}$ and the illumination map $I^{k}$ estimated from $L_{n}^{k-1}$. However, unsupervised removal of severe noise N in low-light light field images through a single denoising process is a challenging task. Therefore, we propose a Content-associated Deep Compensation (CDC) module to mitigate the negative effects of noise and illumination map errors on the optimization process. By leveraging the CDC module’s output $\bm{\delta}^{k}$, we approximate and replace the noise term $\textbf{N}^{k}$ in Eq. (4), thereby improving the robustness and accuracy of the optimization process.
@@ IV-D Deep Optimization
With the estimated illumination map $\textbf{I}^{k}$ and content-associated deep compensation $\bm{\delta}^{k}$ in the $k$-th iteration, we replace the $\textbf{N}^{k}$ in Eq. (4) by $\bm{\delta}^{k}$, and address the optimization problem through a deep optimization module that comprises a defined computation and a deep regularization term. Specifically, we construct the deep implicit regularization term by utilizing pseudo-explicit feature interaction modules with a dense-skip connection. The solution can be expressed as
@@ V-A Experiment Settings
Datasets. Both synthetic LF images from the Inria [53] and HCI [54] benchmarks and real-world LF images captured with a Lytro Illum camera provided by L3F dataset [13] were employed to train and test. Specifically, for the synthetic LF dataset, 22 LF images of dimensions $5\times 5\times 512\times 512$ from HCI and 33 LF images of the same dimensions from Inria were selected to form the training set. The test set contains 2 LF images from HCI and 4 LF images from Inria. The data simulation process follows the methodology described in [39], employing a low-light factor of $0.2$ and a Gaussian noise deviation of $0.25$. For the L3F dataset [13], we use the same training and test set partitioning provided by the dataset, and the dimensions of the LF data are $7\times 7\times 434\times 625$. Note that L3Fnet needs additional SAIs as the input to their view reconstruction block, so the input angular resolution of L3FNet for synthetic LF and real-world LF are $7\times 7$ and $9\times 9$, respectively.
Figure 9: Visual comparisons of enhanced center SAI by different methods on the L3F dataset [13]. Selected regions have been zoomed in for better comparison. The EPI and zoomed selected regions of the input image are the original images ×10,20,30$\times$ 10,20,30× 10 , 20 , 30 for L3F-20, L3F-50, and L3F-100, respectively. It is recommended to view this figure by zooming in.
@@ V-B Comparison with State-of-the-Art Methods
Quantitative comparisons on the synthetic LF dataset. Two strategies were used to simulate the low-light LF images. One is to follow [13] to simulate the data in a fixed illumination environment and use the same photoelectron amplification factor $\alpha=0.2$, denoted as Syn-F. The other is that the simulation data is in a dynamic illumination environment, and the photoelectron amplification factor changes within a particular range $\alpha\in\left[0.1,0.3\right]$, denoted as Syn-D. The standard variance of the adding zero-mean Gaussian noise is 20 and 15, respectively.
@@
Visual comparisons. Fig. 8 and Fig. 9 show the visual comparisons of enhanced LF images by different methods. It can be observed that our method produces better visual results than all the compared methods under all tasks. Specifically, our method can produce sharper edges at the occlusion boundaries and better high-frequency details at texture regions than other approaches. Moreover, the visual effects and overall brightness of our results are closest to the ground truth. As shown in Fig. 8, Ours and Ours-D can enhance the low-light data with intense noise and obtain clear images. For the LF images with great challenges, such as extremely low light or noise, all the results (including Ours) inevitably suffer from a local color distortion.
Figure 10: Visual comparisons of the center depth maps derived by applying the LF depth estimation method in [4] to the enhanced LF images by different methods. Note that the GT depth of Syn-F is available, while the estimated depth of L3F-20 is derived from the GT image.